# Robust Decision Transformer: Tackling Data Corruption in Offline RL via Sequence Modeling

### v1

@@ Abstract
Learning policies from offline datasets through offline reinforcement learning (RL) holds promise for scaling data-driven decision-making and avoiding unsafe and costly online interactions. However, real-world data collected from sensors or humans often contains noise and errors, posing a significant challenge for existing offline RL methods. Our study indicates that traditional offline RL methods based on temporal difference learning tend to underperform Decision Transformer (DT) under data corruption, especially when the amount of data is limited. This suggests the potential of sequential modeling for tackling data corruption in offline RL. To further unleash the potential of sequence modeling methods, we propose Robust Decision Transformer (RDT) by incorporating several robust techniques. Specifically, we introduce Gaussian weighted learning and iterative data correction to reduce the effect of corrupted data. Additionally, we leverage embedding dropout to enhance the model’s resistance to erroneous inputs. Extensive experiments on MoJoCo, KitChen, and Adroit tasks demonstrate RDT’s superior performance under diverse data corruption compared to previous methods. Moreover, RDT exhibits remarkable robustness in a challenging setting that combines training-time data corruption with testing-time observation perturbations. These results highlight the potential of robust sequence modeling for learning from noisy or corrupted offline datasets, thereby promoting the reliable application of offline RL in real-world tasks.
@@ 1 Introduction
Offline reinforcement learning (RL) aims to derive near-optimal policies from fully offline datasets (Levine et al.,, 2020; Wang et al.,, 2018; Fujimoto et al.,, 2019; Kumar et al.,, 2020), thereby reducing the need for costly and potentially unsafe online interactions with the environment. However, offline RL encounters a significant challenge known as distribution shift (Levine et al.,, 2020), which can lead to performance degradation. To address this challenge, several offline RL algorithms impose policy constraints (Wang et al.,, 2018; Fujimoto et al.,, 2019; Fujimoto and Gu,, 2021; Kostrikov et al.,, 2021; Park et al.,, 2024) or maintain pessimistic values for out-of-distribution (OOD) actions (Kumar et al.,, 2020; An et al.,, 2021; Bai et al.,, 2022; Yang et al., 2022a, ; Ghasemipour et al.,, 2022), ensuring that the learned policy aligns closely with the training distribution. Although the majority of traditional offline RL methods rely on temporal difference learning, an alternative promising paradigm for offline RL emerges in the form of sequence modeling (Chen et al.,, 2021; Janner et al.,, 2021; Furuta et al.,, 2021; Shi et al.,, 2023; Wu et al.,, 2024). Unlike traditional RL methods, Decision Transformer (DT) (Chen et al.,, 2021), a representative sequence modeling method, treats RL as a supervised learning task, predicting actions directly from sequences of reward-to-gos, states, and actions. Prior work (Bhargava et al.,, 2023) suggests that DT excels at handling tasks with sparse rewards and suboptimal quality data, showcasing the potential of sequence modeling for real-world applications.
@@
When deploying offline RL in practical settings, dealing with noisy or corrupted data resulting from data collection or malicious attacks is inevitable (Zhang et al.,, 2020; Liang et al.,, 2024; Yang et al., 2024a, ). Therefore, robustly learning policies from such data becomes crucial for the deployment of offline RL. A series of prior works (Zhang et al.,, 2022; Ye et al., 2024b, ; Wu et al.,, 2022; Chen et al.,, 2024; Ye et al., 2024a, ) focus on the theoretical properties and certification of offline RL under data corruption. Notably, a recent study by Ye et al., 2024b proposes an uncertainty-weighted offline RL algorithm with Q ensembles to address reward and dynamics corruption, while Yang et al., 2024b enhance the robustness of Implicit Q-Learning (IQL) (Kostrikov et al.,, 2021) against corruptions on all elements using Huber loss and quantile Q estimators. These advancements, predominantly based on temporal difference learning without utilizing sequence modeling methods, prompt the intriguing question: Can sequential modeling method effectively handle data corruption in offline RL?
@@
In this study, we conduct a comparative analysis of DT with traditional offline RL methods under different data corruption settings. Surprisingly, we observe that traditional methods relying on temporal difference learning struggle in limited data settings and notably underperform DT. Moreover, DT outperforms prior methods under state attack scenarios. These observations highlight the promising potential of sequential modeling for addressing data corruption challenges in offline RL. To further unlock the capabilities of sequential modeling methods under data corruption, we propose a novel algorithm called Robust Decision Transformer (RDT). RDT incorporates three robust techniques for DT, including Gaussian weighted learning, iterative data correction, and embedding dropout to mitigate the impact of corrupted data.
@@
Through comprehensive experiments conducted on a diverse set of tasks, including MoJoCo, KitChen, and Adroit, we demonstrate that RDT outperforms conventional temporal difference learning and sequence modeling approaches under both random and adversarial data corruption scenarios. Additionally, RDT exhibits remarkable robustness in a challenging setting that combines training-time data corruption with testing-time observation perturbations. Our study emphasizes the significance of robust sequential modeling and offers valuable insights that would contribute to the trustworthy deployment of offline RL in real-world applications.
@@ 2 Preliminaries
2.1 RL and Offline RL
RL is generally formulated as a Markov Decision Process (MDP) defined by a tuple $(S,A,P,r,\gamma)$. This tuple comprises a state space $S$, an action space $A$, a transition function $P$, a reward function $r$, and a discount factor $\gamma\in[0,1)$. The objective of RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative return:
$\max_{\pi}\mathbb{E}_{s_{0}\sim\rho_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right],$
where $\rho_{0}$ denotes the distribution of initial states. In offline RL, the objective is to optimize the RL objective with a previously collected dataset $\mathcal{D}=\{(s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)})_{t=0}^{T-1}\}_{i=0}^{N-1}$. which consists of trajectories of length $T$ and a total of $N$ trajectories. The agent cannot directly interact with the environment during the offline phase.
2.2 Decision Transformer (DT)
Different from the MDP framework, DT models decision making from offline datasets as a sequential modeling problem. The $i$-th trajectory $\tau^{(i)}$ of length $T$ in dataset $\mathcal{D}$ is reorganized into a sequence of return-to-go $R_{t}^{(i)}$, state $s_{t}^{(i)}$, action $a_{t}^{(i)}$:
$\tau^{(i)}=\left(R_{0}^{(i)},s_{0}^{(i)},a_{0}^{(i)},\ldots,R_{T-1}^{(i)},s_{T-1}^{(i)},a_{T-1}^{(i)}\right)$ (1)
@@ 2.2 Decision Transformer (DT)
$\mathcal{L}_{DT}(\theta)=\mathbb{E}_{\tau^{(i)}\sim\mathcal{D}}\left[\frac{1}{K}\sum_{t=0}^{K-1}\|\pi_{\theta}(\tau_{t-K+1:t-1}^{(i)},R_{t}^{(i)},s_{t}^{(i)})-a_{t}^{(i)}\|_{2}^{2}\right]$ (2)
@@
2.3 Data Corruption in Offline RL
In prior works (Ye et al., 2024b, ; Yang et al., 2024b, ), the data is stored in transitions, and data corruption is performed on individual elements (state, action, reward, next-state) of each transition. This approach does not align well with trajectory-based sequential modeling methods like DT. In this paper, we consider a unified trajectory-based storage, where corrupting a next-state in a transition corresponds to corrupting a state in the subsequent transition, while corruption for rewards and actions is consistent with prior works. To elaborate, an original trajectory is denoted as $\tau_{\rm{origin}}^{(i)}=\left(s_{0}^{(i)},a_{0}^{(i)},r_{0}^{(i)},\ldots,s_{T-1}^{(i)},a_{T-1}^{(i)},r_{T-1}^{(i)}\right)$, which can be reorganized into the sequence data of DT in Eq. 1 or split into $T-1$ transitions $\left(s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)}\right)_{t=0}^{T-2}$ for MDP-based methods. Note that here only are three independent elements (i.e., states, actions, and rewards) under the trajectory-based storage formulation.
Data corruption injects random or adversarial noise into the original states, actions, and rewards. Random corruption adds random noise to the affected elements in the datasets, resulting in a corrupted trajectory denoted as $\tau_{\rm{corrupt}}^{(i)}=\left(\hat{s}_{0}^{(i)},\hat{a}_{0}^{(i)},\hat{r}_{0}^{(i)},\ldots,\hat{s}_{T-1}^{(i)},\hat{a}_{T-1}^{(i)},\hat{r}_{T-1}^{(i)}\right)$. For instance, $\hat{s}_{0}^{(i)}=s_{0}^{(i)}+\lambda\cdot\textrm{std}(s),\lambda\sim\rm{Uniform}[-\epsilon,\epsilon]^{d_{s}}$ ($d_{s}$ is the dimensions of state, and $\epsilon$ is the corruption scale) and $\textrm{std}(s)$ is the $d_{s}$-dimensional standard deviation of all states in the offline dataset. In contrast, adversarial corruption uses Projected Gradient Descent attack (Madry et al.,, 2017) with pretrained value functions. More details about data corruption are provided in Appendix C.1.
@@ 3 Sequence Modeling for Offline RL with Data Corruption
We aim to answer the question of whether sequential modeling methods can effectively handle data corruption in offline RL in Section 3.1. To achieve this, we compare DT with prior offline RL methods in the context of data corruption. Based on the insights gained from the motivating example, we then propose enhancements for improving the robustness of DT in Section 3.2.
Figure 1: Average normalized scores of offline RL algorithms under random data corruption across three MoJoCo tasks (Halfcheetah, Walker and Hopper) using "medium-replay-v2" datasets. Many offline RL algorithms experience a substantial performance decline when subjected to data corruption. Notably, DT demonstrates superior robustness in the face of state corruption and limited data regimes.
@@ 3.1 Motivating Example
As illustrated in Figure 1, we compare DT with various offline RL algorithms under data corruption. We apply random corruption introduced in Section 2.3 on states, actions, and rewards. Specifically, we perturb $30\%$ of the transitions in the dataset and introduce random noise at a scale of 1.0 standard deviation. In addition to the conventional full dataset setting employed in previous studies (Ye et al., 2024b, ; Yang et al., 2024b, ), we also explore a scenario with a reduced dataset, comprising only $10\%$ of the original trajectories. Our results reveal that most offline RL methodologies are susceptible to data corruption, particularly in the limited data setting. In contrast, DT exhibits remarkable robustness to data corruption, particularly in cases of state corruption and all limited dataset settings. This resilience can be attributed to DT’s formulation based on sequence modeling and its supervised training paradigm. We will later discuss more in-depth about the important components impacting the robustness of DT in Section D.1. In the full dataset setting, it is worth noting that DT exhibits slightly lower performance compared to RIQL and UWMSG when subjected to action and reward attacks, indicating room for further improvement.
Overall, the results demonstrate DT can outperform prior imitation-based and temporal-difference based methods without any robustness enhancement. This observation raises an intriguing question:
How can we further unleash the potential of sequential modeling in addressing data corruption with limited dataset in offline RL?
@@ 3.2 Robust Decision Transformer
To enhance the robustness of DT against various data corruption, we propose Robust Decision Transformer (RDT) by incorporating three components: embedding dropout (Section 3.2.1), Gaussian weighted learning (Section 3.2.2), and iterative data correction (Section 3.2.3). The overall framework is depicted in Figure 2. Notably, RDT predicts actions and rewards (rather than reward-to-gos) on top of the DT. Given the typically high dimensionality of states, we avoid predicting states to mitigate potential negative impacts. Additionally, rewards provide more direct supervision for the policy compared to reward-to-gos, which are also dependent on future actions.
@@ 3.2.1 Embedding Dropout
In the setting of data corruption, corrupted states, actions, and rewards can supply shifted inputs or erroneous features to the model. This can lead to a performance drop if the model overfits to particular harmful features. Therefore, learning a robust representation is important to enhance the model’s resilience against data corruption scenario. To achieve this, we employ embedding dropout, which encourages the model to learn more robust embedding representations and avoid overfitting (Merity et al.,, 2017).
To clarify, we define three hidden embeddings after the linear projection layer as $h_{R_{t}}$, $h_{s_{t}}$, and $h_{a_{t}}$, represented as:
$h_{R_{t}}=\phi_{R}(R_{t})+\phi_{t}(t),\quad h_{s_{t}}=\phi_{s}(s_{t})+\phi_{t}(t),\quad h_{a_{t}}=\phi_{a}(a_{t})+\phi_{t}(t).$
@@
$\tilde{h}_{R_{t}}=\mathcal{M}(p)\otimes h_{R_{t}},\quad\tilde{h}_{s_{t}}=\mathcal{M}(p)\otimes h_{s_{t}},\quad\tilde{h}_{a_{t}}=\mathcal{M}(p)\otimes h_{a_{t}}.$ (3)
The function $\mathcal{M}(\cdot)$ takes a probability as input and outputs a binary mask to determine the dimension inclusion. We use the same drop probabilities $p$ for different elements embedding.

### v3

# Tackling Data Corruption in Offline Reinforcement Learning via Sequence Modeling

@@ Abstract
Learning policy from offline datasets through offline reinforcement learning (RL) holds promise for scaling data-driven decision-making while avoiding unsafe and costly online interactions. However, real-world data collected from sensors or humans often contains noise and errors, posing a significant challenge for existing offline RL methods, particularly when the real-world data is limited. Our study reveals that prior research focusing on adapting predominant offline RL methods based on temporal difference learning still falls short under data corruption when the dataset is limited. In contrast, we discover that vanilla sequence modeling methods, such as Decision Transformer, exhibit robustness against data corruption, even without specialized modifications. To unlock the full potential of sequence modeling, we propose Robust Decision Transformer (RDT) by incorporating three simple yet effective robust techniques: embedding dropout to improve the model’s robustness against erroneous inputs, Gaussian weighted learning to mitigate the effects of corrupted labels, and iterative data correction to eliminate corrupted data from the source. Extensive experiments on MuJoCo, Kitchen, and Adroit tasks demonstrate RDT’s superior performance under various data corruption scenarios compared to prior methods. Furthermore, RDT exhibits remarkable robustness in a more challenging setting that combines training-time data corruption with test-time observation perturbations. These results highlight the potential of sequence modeling for learning from noisy or corrupted offline datasets, thereby promoting the reliable application of offline RL in real-world scenarios. Our code is available at https://github.com/jiawei415/RobustDecisionTransformer.
@@ 1 Introduction
Offline reinforcement learning (RL) aims to derive near-optimal policies from fully offline datasets (Levine et al., 2020; Kumar et al., 2020; Fujimoto et al., 2019; Wang et al., 2018), thereby reducing the need for costly and potentially unsafe online interactions with the environment. However, offline RL encounters a significant challenge known as distribution shift (Levine et al., 2020), which can lead to performance degradation. To address this challenge, several offline RL algorithms impose policy constraints (Wang et al., 2018; Fujimoto et al., 2019; Fujimoto & Gu, 2021; Kostrikov et al., 2021; Park et al., 2024) or maintain pessimistic values for out-of-distribution (OOD) actions (Kumar et al., 2020; An et al., 2021; Bai et al., 2022; Yang et al., 2022a; Ghasemipour et al., 2022), ensuring that the learned policy aligns closely with the training distribution. Although the majority of traditional offline RL methods rely on temporal difference learning, an alternative promising paradigm for offline RL emerges in the form of sequence modeling (Chen et al., 2021; Janner et al., 2021; Shi et al., 2023; Wu et al., 2024). Unlike traditional RL methods, Decision Transformer (DT) (Chen et al., 2021), a representative sequence modeling method, treats offline RL as a supervised learning task, predicting actions directly from sequences of reward-to-gos, states, and actions. Prior work (Bhargava et al., 2023) suggests that DT excels at handling tasks with sparse rewards and suboptimal quality data, showcasing the potential of sequence modeling for real-world applications.
@@
When deploying offline RL in practical settings, dealing with noisy or corrupted data resulting from data collection or malicious attacks is inevitable (Zhang et al., 2020; 2021; Liang et al., 2024; Yang et al., 2024a). Consequently, robust policy learning from such data is essential for the successful deployment of offline RL. A series of prior works (Zhang et al., 2022; Ye et al., 2024b; Wu et al., 2022; Chen et al., 2024; Ye et al., 2024a) focus on the theoretical properties and certification of offline RL under data corruption. Notably, Ye et al. (2024b) propose an uncertainty-weighted offline RL algorithm with Q ensembles to address reward and dynamics corruption, while Yang et al. (2024b) enhance the robustness of Implicit Q-Learning (IQL) (Kostrikov et al., 2021) against corruptions on all elements using Huber loss and quantile Q estimators. These advancements have primarily concentrated on temporal difference learning, dedicating significant effort to learning robust Q functions from corrupted data. This approach brings up an intriguing question: by bypassing the Q function, can sequence modeling methods effectively handle data corruption in offline RL?
@@
In this study, we conduct a comparative analysis between prior robust offline RL methods and DT (without any robust modifications) under various data corruption settings. Surprisingly, we find that traditional methods relying on temporal difference learning struggle in settings with limited data and significantly underperform compared to DT. Additionally, DT demonstrates superior performance over previous methods in scenarios involving state attacks. These findings highlight the promising potential of sequence modeling for addressing data corruption challenges in offline RL. To further unlock the capabilities of sequence modeling methods under data corruption, inspired by successes in the supervised learning domain (Song et al., 2022; Chang et al., 2017; Reed et al., 2014; Gal & Ghahramani, 2016), we propose Robust Decision Transformer (RDT). RDT aims to boost the robustness of DT in three aspects: model structure, loss function, and data refinement. To achieve this, RDT incorporates three simple yet highly effective techniques to mitigate the impact of corrupted data: embedding dropout, Gaussian weighted learning, and iterative data correction. These simple modifications are designed to address the fundamental limitations of the standard DT and unlock the potential of sequence modeling for data corruption.
@@
To comprehensively study RDT’s robustness under data corruption with limited data, we constructed a variety of small datasets from MuJoCo, Kitchen, and Adroit, creating a challenging benchmark for current robust offline RL methods. Through our experiments, we demonstrate that RDT outperforms conventional temporal difference learning and sequence modeling approaches under both random and adversarial data corruption scenarios, achieving a substantial 29% improvement in overall performance over DT. Furthermore, we show that after learning from corrupted data, RDT exhibits remarkable robustness against testing-time observation perturbations, indicating its potential to address both training-time and test-time attacks simultaneously. Our study emphasizes the significance of robust sequence modeling and offers valuable insights for the trustworthy deployment of offline RL in real-world applications.
@@ 2 Preliminaries
RL and Offline RL.
RL is generally formulated as a Markov Decision Process (MDP) defined by a tuple $(S,A,P,r,\gamma)$. This tuple comprises a state space $S$, an action space $A$, a transition function $P$, a reward function $r$, and a discount factor $\gamma\in[0,1]$. The objective of RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative return: $\max_{\pi}\mathbb{E}_{s_{0}\sim\rho_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{T-1}\gamma^{t}r(s_{t},a_{t})\right],$ where $\rho_{0}$ denotes the distribution of initial states and $T$ is the trajectory length. In offline RL, the objective is to optimize the RL objective with a previously collected dataset $\mathcal{D}=\{(s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)})_{t=0}^{T-1}\}_{i=0}^{N-1}$, which contains a total of $N$ trajectories. The agent cannot directly interact with the environment during the offline phase.
Decision Transformer (DT).
DT models decision-making from offline datasets as a sequence modeling problem. The $i$-th trajectory $\tau^{(i)}$ of length $T$ in dataset $\mathcal{D}$ is reorganized into a sequence of return-to-go $R_{t}^{(i)}$, state $s_{t}^{(i)}$, action $a_{t}^{(i)}$:
$\tau^{(i)}=\left(R_{0}^{(i)},s_{0}^{(i)},a_{0}^{(i)},\ldots,R_{T-1}^{(i)},s_{T-1}^{(i)},a_{T-1}^{(i)}\right)\,.$ (1)
@@
$\mathcal{L}_{DT}(\theta)=\mathbb{E}_{\tau^{(i)}\sim\mathcal{D}}\left[\frac{1}{K}\sum_{t=0}^{K-1}\|\pi_{\theta}(\tau_{t-K+1:t-1}^{(i)},R_{t}^{(i)},s_{t}^{(i)})-a_{t}^{(i)}\|_{2}^{2}\right]\,,$ (2)
@@
Data Corruption in Offline RL.
In prior works (Ye et al., 2024b; Yang et al., 2024b), the data is stored in transitions, and data corruption is performed on individual elements (state, action, reward, next-state) of each transition. This approach does not align well with trajectory-based sequence modeling methods like DT. In this paper, we consider a unified trajectory-based storage, where corrupting a next-state in a transition corresponds to corrupting a state in the subsequent transition, while corruption for rewards and actions is consistent with prior works. To elaborate, an original trajectory is denoted as $\tau_{\rm{origin}}^{(i)}=\left(s_{0}^{(i)},a_{0}^{(i)},r_{0}^{(i)},\ldots,s_{T-1}^{(i)},a_{T-1}^{(i)},r_{T-1}^{(i)}\right)$, which can be reorganized into the sequence data of DT in Eq. 1 or split into $T-1$ transitions $\left(s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)}\right)_{t=0}^{T-2}$ for MDP-based methods. Note that here are only three independent elements (i.e., states, actions, and rewards) under the trajectory-based storage formulation.
Data corruption injects random or adversarial noise into the original states, actions, and rewards. Random corruption adds random noise to the affected elements in the datasets, resulting in a corrupted trajectory denoted as $\tau_{\rm{corrupt}}^{(i)}=\left(\hat{s}_{0}^{(i)},\hat{a}_{0}^{(i)},\hat{r}_{0}^{(i)},\ldots,\hat{s}_{T-1}^{(i)},\hat{a}_{T-1}^{(i)},\hat{r}_{T-1}^{(i)}\right)$. For instance, $\hat{s}_{0}^{(i)}=s_{0}^{(i)}+\lambda\cdot\textrm{std}(s),\lambda\sim\rm{Uniform}[-\epsilon,\epsilon]^{d_{s}}$ ($d_{s}$ is the dimensions of state, and $\epsilon$ is the corruption scale) and $\textrm{std}(s)$ is the $d_{s}$-dimensional standard deviation of all states in the offline dataset. In contrast, adversarial corruption uses Projected Gradient Descent attack (Madry et al., 2017) with pretrained value functions. Specifically, we introduce learnable noise to the states or actions and then optimize this noise by minimizing the pretrained value functions through gradient descent. More details about data corruption are provided in Appendix C.1.
@@ 3 Sequence Modeling for Offline RL with Data Corruption
We aim to answer the question of whether sequence modeling methods can effectively handle data corruption in offline RL in Section 3.1. To achieve this, we compare DT with prior offline RL methods in the context of data corruption. Based on the insights gained from the motivating example, we then propose enhancements for improving the robustness of DT in Section 3.2.
Figure 1: Average normalized scores of offline RL algorithms under random data corruption across three MuJoCo tasks (halfcheetah, walker2d, and hopper) using “medium-replay-v2” datasets. Many offline RL algorithms experience substantial performance declines when subjected to data corruption. In contrast, DT demonstrated remarkable robustness, particularly in the $10\%$ data regime.
@@ 3.1 Motivating Example
As illustrated in Figure 1, we compare DT with various offline RL algorithms under data corruption. We apply random corruption introduced in Section 2 on states, actions, and rewards. Specifically, we perturb $30\%$ of the transitions in the dataset and introduce random noise at a scale of 1.0 standard deviation. In addition to the conventional full dataset setting employed in previous studies (Ye et al., 2024b; Yang et al., 2024b), we also explore a scenario with a reduced dataset, comprising only $10\%$ of the original trajectories. Our results reveal that most offline RL methods are highly vulnerable to data corruption, particularly in more challenging, limited data settings. In contrast, DT exhibits remarkable robustness to data corruption, especially in scenarios involving state corruption and limited dataset conditions. This resilience can be attributed to DT’s formulation based on sequence modeling and its supervised training paradigm. We will delve deeper into the critical components affecting DT’s robustness in Section D.1.
Overall, the results demonstrate DT can outperform prior imitation-based and temporal difference based methods without any robustness enhancement. This observation raises an interesting question:
How can we further unleash the potential of sequence modeling in addressing data corruption with limited dataset in offline RL?
@@ 3.2 Robust Decision Transformer
To improve the robustness of DT against various data corruptions, we draw inspiration from successes in the supervised learning domain (Song et al., 2022; Chang et al., 2017; Reed et al., 2014; Gal & Ghahramani, 2016) to introduce the Robust Decision Transformer (RDT). RDT enhances DT across three key aspects: model structure, loss function, and data refinement. This is achieved through the incorporation of three simple yet effective components: embedding dropout (Section 3.2.1), Gaussian weighted learning (Section 3.2.2), and iterative data correction (Section 3.2.3). These simple modifications are designed to tackle the fundamental limitations of the standard DT and help unlock the potential of sequence modeling for data corruption. The overall framework is illustrated in Figure 2. Notably, RDT predicts both actions and rewards (instead of reward-to-gos). Given the typically high dimensionality of states, we avoid predicting states to mitigate potential negative impacts as discussed in Section D.14. Additionally, rewards provide more direct supervision for the policy compared to reward-to-gos, which also depend on future actions.
@@ 3.2.1 Embedding Dropout
In the context of data corruption, corrupted states, actions, and rewards can introduce shifted inputs or erroneous features to the model. This can lead to a performance drop if the model overfits these harmful features. Therefore, developing a robust representation is crucial to enhance the model’s resilience against data corruption.
Several studies (Carroll et al., 2022; Hu et al., 2023) utilize element-wise masking to learn robust representations by randomly masking state, return-to-go, or action elements in a trajectory, leveraging the redundancy in the trajectory’s information. However, we discovered that corrupted trajectories contain less redundant information. Discarding raw elements directly can lead to notable information loss, ultimately resulting in decreased performance (see Figure 3(a)). To address this, we propose randomly dropping dimensions in the feature space rather than masking raw elements. Specifically, we employ embedding dropout, which encourages the model to learn more robust embedding representations while preventing overfitting (Merity et al., 2017).
To clarify, we define three hidden embeddings after the linear projection layer as $h_{R_{t}}$, $h_{s_{t}}$, and $h_{a_{t}}$, formulated as:
$h_{R_{t}}=\phi_{R}(R_{t})+\phi_{t}(t),\quad h_{s_{t}}=\phi_{s}(s_{t})+\phi_{t}(t),\quad h_{a_{t}}=\phi_{a}(a_{t})+\phi_{t}(t).$ (3)
@@
$\tilde{h}_{R_{t}}=\mathcal{M}(p)\odot h_{R_{t}},\quad\tilde{h}_{s_{t}}=\mathcal{M}(p)\odot h_{s_{t}},\quad\tilde{h}_{a_{t}}=\mathcal{M}(p)\odot h_{a_{t}}.$ (4)
The function $\mathcal{M}(\cdot)$ takes a probability $p\in[0,1]$ as input and outputs a binary mask with the same shape as embeddings to determine the dimension inclusion. We apply the same drop probabilities $p$ within a moderate range $(0.1,0.3)$ for different element embeddings.
Figure 3: (a) Comparing dropout methods under state attack: Embedding dropout outperforms directly dropping the entire state (DeFog) or dropping dimensions on the raw state. (b) Gaussian weighted learning under action attack: Gaussian weighted learning (DT w. GWL) alleviates overfitting to the corrupted data and slightly minimizes the loss on clean data. (c) Iterative data correction (DT w. IDC ) under action attack: The MSE between corrected and oracle data gradually decreases to near zero.